Herman dries 60 mph for one hour and 80 mph for the next hour. What is his average speed?
Answer by jim_thompson5910(35256)

"Herman dries 60 mph for one hour" so he drove 60*1 = 60 miles so far

Then drives "80 mph for the next hour" which means he drove another 80*1 = 80 miles

In total, 60+80 = 140 miles have been driven over 1+1 = 2 hours.

Average Speed = (140 miles)/(2 hours) = 70 miles per hour